OpenStudy (anonymous):

Evaluate the expression. 49^(3/2)

OpenStudy (anonymous):

i got 8 but it was wrong

OpenStudy (anonymous):

lol i answered the wrong problem :P

OpenStudy (anonymous):

\[49^{3/2}=\sqrt{{49^{3}}}\]

OpenStudy (anonymous):

yea i see that. but i got some huge number.w hat do i do then?

OpenStudy (anonymous):

\[\sqrt{49^3}=\sqrt{49\times49\times49}=\sqrt{49}\times \sqrt{49}\times \sqrt{49}=7^3\]\[7^3 = 7\times7\times7=49\times7=343\]

OpenStudy (anonymous):

ohhh. i was doing 49*49*49.

OpenStudy (anonymous):

You can do that, but don't forget to take the square root of 49*49*49.

OpenStudy (anonymous):

yea, i got some 6 digit number and couldn't figure out what to do. thanks!

OpenStudy (anonymous):

No problem!